OpenStudy (something):

Amir pitched a baseball at an initial height of 5.9 feet with a velocity of 73.16 feet per second. About how long will it take the ball Amir pitched to hit the ground?

OpenStudy (anonymous):

is gravity taken into account?

OpenStudy (something):

i dont think so...

OpenStudy (anonymous):

v = s/t 73.16 = 5.9/t t=0.08

OpenStudy (amistre64):

it depends on the angle at which it left the pitch

OpenStudy (anonymous):

yes, don't think he threw the ball straight down at the ground